OpenStudy (anonymous):

Square root 64 Y Squared

OpenStudy (anonymous):

is this \[\sqrt{64y^2}\]or\[\sqrt{64y}^2\]or\[\sqrt{64}y^2\]?

OpenStudy (anonymous):

yes the first one

OpenStudy (anonymous):

sqrt(64y^2) is the answer

OpenStudy (anonymous):

\[\sqrt{64y^2}=\sqrt{64}\times\sqrt{y^2}\]

OpenStudy (anonymous):

what's the square root of 64?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

what's the square root of y^2?

OpenStudy (anonymous):

wouldn't it just be y?

OpenStudy (anonymous):

it would be y for all y>0, or -y for all y<0

OpenStudy (anonymous):

in other words, |y|

OpenStudy (anonymous):

ohhh I see! Thanks

OpenStudy (anonymous):

you're welcome